# Towards a Formal Theory of the Need for Competence via Computational Intrinsic Motivation

### v1

@@ Introduction
Given an agent has set itself a goal (task), if an agent has a mechanism to estimate its progress towards a chosen goal, it can use that estimate as a proxy for its own performance on the task. An example system of this kind is Reinforcement Learning With Imagined Goals (RIG) \@BBOPcitep\@BAP\@BBN(Nair et al.,, 2018)\@BBCP. RIG encodes the agent’s observations into a lower-dimensional latent space from which the agent samples goals to pursue. Defining the reward as the negative distance between the goal and the current state in latent space encourages the agent to move its observations closer to the goal:
@@
In contrast to DIAYN, in which the agent learns to master a fixed number of skills, in RIG, the agent selects from an infinite number of goals (the space is continuous). This results in a sort of smoothness of the space that the agent can benefit from: if one goal is sampled near another that the agent has already learned to achieve, it may be able to generalise and use some parts of an existing policy to more efficiently learn to achieve the new goal. In the RIG system, goals take on the intuitive meaning as desired states for the agent to reach (discussed in Background: Computational IM & RL), and “performing well” at an intended task is operationalised as minimising the semantic distance between the agent’s observations and its self-generated goals. Over time, RIG agents learn to generate goals similar to the observations they make of their environment, while simultaneously learning the skills necessary to consistently perform well (reach the goals).
@@
For example, in the Continual Universal Reinforcement learning with Intrinsically mOtivated sUbstitutionS (CURIOUS) system by \@BBOPcitet\@BAP\@BBNColas et al., (2019)\@BBCP, instead of considering task performance as distance from the goal, the system considers its capability to achieve its goals. Specifically, the agent is motivated to pursue goals in certain parts of the sensory space in which it believes to be increasingly or decreasingly successful at achieving its goals, estimated by measuring how successful it has been in the recent past—that is, its LP. The underpinning logic of the system is that such goals will be of optimal difficulty, neither too easy nor too difficult. In the CURIOUS experiments, an agent manipulates cubes by learning to control a robotic arm (\@BBOPcitep\@BAP\@BBNColas et al.,, 2019, Supplementary p. 3\@BBCP). Each observation includes information about the gripper and the objects of interest (e.g., their position and rotation), and each goal is a desired observation. The high-dimensional observation space makes learning challenging. Furthermore, some parts of the sensory space are irrelevant for achieving certain goals—for example, if the goal is to manipulate one of several cubes, the agent might not need information about the other cubes. To ignore such irrelevant information in its formulation of goals, CURIOUS modularises the observation space, which is also the goal space, with each module being a specific, hand-defined subspace of the space. A CURIOUS agent estimates its own LP for each module, and uses the absolute values of the LP estimates to compute a probability distribution for selecting the module to generate a goal within \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2019, p. 4)\@BBCP. This mechanism helps the agent avoid modules containing goals that are too difficult or for which it already has sufficient skill, as well as to return to skills it is forgetting \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2019, pp. 2–3)\@BBCP. In this way, the selection of different goals is driven by self-assessment of task performance.
The subjective LP, for a given module, $LP(M_{i})$, is defined as the derivative of subjective competence, $C(M_{i})$, with respect to time, where $C(M_{i})$ is defined as the mean over results from self-evaluated trials performed by the agent (see \@BBOPcitep\@BAP\@BBNColas et al.,, 2019, p. 4\@BBCP). The results are stored in a queue of preset length, and each result is binary: either 1 for a successful trial or 0 for a failure. This can be viewed as the frequentist estimate of the probability of success. Finally, the probability of choosing a module is computed as:
@@
In the same vein, an emerging ACL paradigm, Unsupervised Environment Design (UED) (for a review see \@BBOPcitep\@BAP\@BBNRutherford et al.,, 2024\@BBCP), formalises the problem of automatically generating tasks for agents based on some criterion, for instance, the difference between the current agent and (in practice, an approximation of) an optimal agent \@BBOPcitep\@BAP\@BBN(Rutherford et al.,, 2024, p. 1)\@BBCP. In UED, environments can be thought of as goals for an agent to pursue (each associated with their own reward function). As part of one such system, Open-endedness via Models of human Notions of Interestingness with Environments Programmed in Code (OMNI-EPIC), agents self-generate novel, interesting, and solvable goals, matched for an agent’s current capabilities \@BBOPcitep\@BAP\@BBN(Faldor et al.,, 2024, p. 2)\@BBCP.
While a significant proportion of ACL methods optimise task difficulty or competence progress with respect to current capabilities, ACL can also motivate agents to optimise goal selection based on other criteria that can be interpreted as proxies for “skill level” (from C3.b). \@BBOPcitet\@BAP\@BBNPong et al., (2020)\@BBCP extend RIG (see C3.a) by weighting the selection of previously visited states used in learning the goal-generation model, encouraging the agent to generate goals resembling less frequently visited states. This weighting can be seen as a heuristic for attending to tasks that are expected to be non-trivial for the agent. As a second example, \@BBOPcitet\@BAP\@BBNLintunen et al., 2024a \@BBCP introduce a method motivating agents to select goals based on how much the agent believes pursuing them will diversify its entire set of skills. As ACL systems often consider the level of skill needed to perform well on the task (where the task could defined as learning a diverse set of skills), autocurricula methods offer a good starting point for formalising C3.b.
@@
An example of a system that explores both C4.a and C4.b—both repeatedly engaging in skills and extending the agent’s skill repertoire—is Intrinsically Motivated Reinforcement Learning (IMRL) described by \@BBOPcitet\@BAP\@BBNBarto et al., (2004)\@BBCP and \@BBOPcitet\@BAP\@BBNSingh et al., (2004)\@BBCP. IMRL extends the agent’s repertoire of skills (C4.b): each time a “salient” event is observed, the agent adds a new skill associated with the salient event.333\@BBOPcitet\@BAP\@BBNBarto et al., (2004)\@BBCP leave to future work what salient means and how to determine whether an event is salient or not. They simply hard-code events like turning on a light or ringing a bell as “salient.” After this first observation, the agent learns how to return to observe the salient event again from any other situation via experience over time—this is the new skill associated with the salient event. Strictly speaking, the skill repertoire expands as a matter of course, not because the agent is intrinsically rewarded for expanding it.
@@
The agent only receives a non-zero intrinsic reward when it reaches a salient state. The intrinsic reward here can be thought of as modelling how surprising it is that the agent reached ${\mathbf{s}}_{t+1}$ immediately from state ${\mathbf{s}}_{t}$—which it has just done—by following the skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$. If the agent expected, without uncertainty, to transition directly from ${\mathbf{s}}_{t}$ to ${\mathbf{s}}_{t+1}$, then $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ is 1 and the reward is 0 (the transition is very unsurprising). If the agent did not expect it could transition from ${\mathbf{s}}_{t}$ to ${\mathbf{s}}_{t+1}$ by following skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$, then $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ is 0 and the reward is 1 (maximally surprising). If the agent is uncertain, $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ will be somewhere in between. However, if the agent expected that it could reach ${\mathbf{s}}_{t+1}$ from ${\mathbf{s}}_{t}$ by following skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$, just that it would take more than a single step, $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ will be discounted based on the length of the path; this discounting is part of how the multi-time model is constructed.
@@
Another example of C4.b, that is, being motivated by growing the number of skills in the agent’s repertoire, is the aforementioned VIC (see C2.a). The VIC reward function motivates the agent to grow its number of effective skills: in expectation, the second term of the reward function (Eq. 2, $\beta$) increases as the distribution over skills becomes more uniform. In theory, this encourages the distribution over skills not to collapse to a small number of skills. In practice, since there are numerous other factors that affect how agents learn, this does not always happen (for empirical results see \@BBOPcitep\@BAP\@BBNEysenbach et al.,, 2019, Appendix E, pp. 17–18\@BBCP).
Advancing the Theory of Competence in SDT
@@
Why should human motivation researchers care about the computational models we have identified? By identifying distinct ways the different facets of the need for competence in SDT (Table 1) can be formalised, we better understand the ways the theory is underspecified. For instance, LP-based methods (discussed in C2, C3, C4) can exemplify skill use as an opportunity to use a particular skill (C2.a), task performance that requires a certain skill or skill level (C3.b), and capacity growth in the sense of observing gains in the strength of one’s skills (C4.a). Given that SDT does not acknowledge distinct facets, finding such a formalism that fits multiple facets supports the theory. On the other hand, some facets are formally distinct. For example, it is unclear how effectance (C1), being motivated by extending one’s repertoire of skills (C4.b), or being motivated by observing that one performs well at an intended task (C3.a) relate to each other and other facets. Different interpretations of the theory might lead to drastically different operationalisations of the need for competence. Our research can alleviate this problem by uncovering ways the theory conflates formally distinct concepts, thus prompting review of the theory.
@@
Another reason human motivation researchers should care about this work is that the computational formalisms reveal underlying preconditions that SDT fails to make explicit, such as that a competence-motivated agent is able to recognise when actions cause effects (C1) and differentiate between distinct skills (C2, C3, C4). Similarly, concepts from intrinsically motivated RL, such as novelty, diversity, and progress towards goals, underlie competence-driven learning in computational contexts, yet are often overlooked in SDT. Our research shines light on mechanisms used to computationally fulfil such preconditions, and these could be investigated for further insights. Additionally, our work can support a cycle of theory development by inspiring new computational models formalising aspects of the theory, which can then be tested empirically to refine the theory.
@@
For computational IM to be effective in informing theory on motivation, more empirical work is needed in the comparison and evaluation of computational IM. While some work exists comparing behaviour induced by different intrinsic reward functions (e.g., \@BBOPcitep\@BAP\@BBNSantucci et al.,, 2012; Biehl et al.,, 2018; Linke et al.,, 2020\@BBCP), as does work studying variables such as the choice of representation (e.g., \@BBOPcitep\@BAP\@BBNBurda et al.,, 2019\@BBCP), more research is needed on the properties of different models, such as how they behave in varying environments and bodies. Showing that RL formalisms can be aligned with SDT uncovers a wealth of resources for motivation research, such as existing implementations of models, simulators, and evaluation metrics. These resources can enable large-scale simulation studies to explore psychologically plausible computational IMs and how they may explain human data, thus contributing to formal competence-related theory. However, we need work to decide how to effectively compare artificial and biological agents (a project partially undertaken by researchers in representational alignment; e.g., \@BBOPcitep\@BAP\@BBNSucholutsky et al.,, 2024\@BBCP).
Further research is also required to understand the extent to which computational IM research intersects with intrinsically motivated learning as understood in psychology. Early work on computational IM emerged at the intersection of developmental psychology and robotics, attempting to model intrinsically motivated human behaviours, but a large portion of state-of-the-art intrinsically motivated RL is not necessarily aligned with relevant goals of understanding the human mind or developing safe (human-aligned) AI systems.
@@
As most formalisms we identified relate to the pursuit of goals, we need to better understand how humans represent and generate goals \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2024, p. 1)\@BBCP. The topic has not been well studied for various reasons, including computational intractability \@BBOPcitep\@BAP\@BBN(Byers et al.,, 2024, p. 1)\@BBCP. Excitingly, goals have recently emerged as a multidisciplinary area of research, cutting across AI, cognitive science, and developmental psychology (e.g., \@BBOPcitep\@BAP\@BBNColas et al.,, 2022; Molinaro and Collins,, 2023; Chu et al.,, 2024; Davidson et al.,, 2024\@BBCP). Recent work by \@BBOPcitet\@BAP\@BBNMolinaro and Collins, (2023, p. 1150)\@BBCP suggests that goals affect many agent dynamics, such as representing the environment, choosing relevant actions, and evaluating rewards.
@@
In summary, we build on recent research arguing how the need for competence is at present poorly specified in SDT. We do so by identifying computational formalisms to inspire empirical work to resolve this issue. We show that the four distinct facets of competence—effectance, skill use, task performance, and capacity growth—can be aligned with existing formalisms from intrinsically motivated RL, highlighting the potential of computational models to provide deeper insights into competence-motivated behaviours. While this is a promising first step, further research is required to empirically study these computational models, inviting collaboration from motivation researchers across disciplines.

### v2

@@ Introduction
Given an agent has set itself a goal (task), if an agent has a mechanism to estimate its progress towards a chosen goal, it can use that estimate as a proxy to track its own performance on the task. An example system of this kind is Reinforcement Learning With Imagined Goals (RIG) \@BBOPcitep\@BAP\@BBN(Nair et al.,, 2018)\@BBCP. RIG encodes the agent’s observations into a lower-dimensional latent space from which the agent samples goals to pursue. Defining the reward as the negative distance between the goal and the current state in latent space encourages the agent to move its observations closer to the goal:
@@
In contrast to DIAYN, in which the agent learns to master a fixed number of skills, in RIG, the agent selects from an infinite number of goals (the space is continuous). This results in a sort of smoothness of the space that the agent can benefit from: if one goal is sampled near another that the agent has already learned to achieve, it may be able to generalise and use some parts of an existing policy to more efficiently learn to achieve the new goal. In the RIG system, goals take on the intuitive meaning as desired states for the agent to reach (discussed in Background: Computational IM & RL), and “performing well” at an intended task is operationalised as minimising the semantic distance between the agent’s observations and its self-generated goals. Over time, RIG agents learn to generate goals that approximate actual observations they make of their environment, while simultaneously learning the skills necessary to consistently perform well (reach the goals).
@@
For example, in the Continual Universal Reinforcement learning with Intrinsically mOtivated sUbstitutionS (CURIOUS) system by \@BBOPcitet\@BAP\@BBNColas et al., (2019)\@BBCP, the system considers its capability to achieve its goals. Specifically, the agent is motivated to pursue goals in certain parts of the sensory space in which it believes itself to be increasingly or decreasingly successful at achieving its goals, estimated by measuring how successful it has been in the recent past—that is, its LP. The underpinning logic of the system is that such goals will be of optimal difficulty, neither too easy nor too difficult. In the CURIOUS experiments, an agent manipulates cubes by learning to control a robotic arm (\@BBOPcitep\@BAP\@BBNColas et al.,, 2019, Supplementary p. 3\@BBCP). Each observation includes information about the gripper and the objects of interest (e.g., their position and rotation), and each goal is a desired observation. The high-dimensional observation space makes learning challenging. Furthermore, some parts of the sensory space are irrelevant for achieving certain goals—for example, if the goal is to manipulate only one of several cubes, the agent might not need information about the other cubes. To ignore such irrelevant information in its formulation of goals, CURIOUS modularises the observation space, which is also the goal space, with each module being a specific, hand-defined subspace of this space. A CURIOUS agent estimates its own subjective LP for each module, and uses the absolute values of its LP estimates to compute a probability distribution for selecting the module to generate a goal within \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2019, p. 4)\@BBCP. This mechanism helps the agent avoid modules containing goals that are too difficult or for which it already has sufficient skill, as well as to return to skills it is forgetting \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2019, pp. 2–3)\@BBCP. In this way, the selection of different goals is driven by self-assessment of task performance.
The subjective LP for a given module, $LP(M_{i})$, is defined as the derivative of subjective competence, $C(M_{i})$, with respect to time, where $C(M_{i})$ is defined as the mean over results from self-evaluated trials performed by the agent (see \@BBOPcitep\@BAP\@BBNColas et al.,, 2019, p. 4\@BBCP). The results are stored in a queue of preset length, and each result is binary: either 1 for a successful trial or 0 for a failure. This can be viewed as the frequentist estimate of the probability of success. Finally, the probability of choosing a module is computed as:
@@
In the same vein, an emerging ACL paradigm, Unsupervised Environment Design (UED) (for a review see \@BBOPcitep\@BAP\@BBNRutherford et al.,, 2024\@BBCP), formalises the problem of automatically generating environments for agents based on some criterion, for instance, the difference between the current agent and (in practice, an approximation of) an optimal agent \@BBOPcitep\@BAP\@BBN(Rutherford et al.,, 2024, p. 1)\@BBCP. In UED, environments with their own reward functions can be thought of as goals for an agent to pursue. As part of one such system, Open-endedness via Models of human Notions of Interestingness with Environments Programmed in Code (OMNI-EPIC), agents self-generate novel, interesting, and solvable goals, appropriate for an agent’s current capabilities \@BBOPcitep\@BAP\@BBN(Faldor et al.,, 2024, p. 2)\@BBCP.
While a significant proportion of ACL methods optimise task difficulty or competence progress with respect to current capabilities, ACL can also motivate agents to optimise goal selection based on other criteria that can be interpreted as proxies for “skill level” (from C3.b). \@BBOPcitet\@BAP\@BBNPong et al., (2020)\@BBCP extended RIG (introduced in C3.a) by weighting the selection of previously visited states used in learning the goal-generation model, encouraging the agent to generate goals resembling less frequently visited states. This weighting can be seen as a heuristic for attending to tasks that are expected to be non-trivial for the agent. As a second example, \@BBOPcitet\@BAP\@BBNLintunen et al., (2024)\@BBCP introduce a method motivating agents to select goals based on how much the agent believes pursuing them will diversify its entire set of skills. As ACL systems often consider the level of skill needed to perform well on the task, they offer a good starting point for formalising C3.b.
@@
An example of a system that explores both C4.a and C4.b—both repeatedly engaging in skills that need strengthening and extending the agent’s skill repertoire—is Intrinsically Motivated Reinforcement Learning (IMRL) described by \@BBOPcitet\@BAP\@BBNBarto et al., (2004)\@BBCP and \@BBOPcitet\@BAP\@BBNSingh et al., (2004)\@BBCP. IMRL extends the agent’s repertoire of skills (C4.b): each time a “salient” event is observed, the agent adds a new skill associated with the salient event.444\@BBOPcitet\@BAP\@BBNBarto et al., (2004)\@BBCP leave to future work what “salient” means and how to determine whether an event is salient or not. They simply hard-code events like turning on a light or ringing a bell as salient. After this first observation, the agent learns how to return to observe the salient event again from any other situation via experience over time—this is the new skill associated with the salient event. Strictly speaking, the skill repertoire expands as a matter of course, not because the agent is intrinsically rewarded for expanding it.
@@
The agent only receives a non-zero intrinsic reward when it reaches a salient state. The intrinsic reward here can be thought of as modelling how “surprising” \@BBOPcitep\@BAP\@BBN(Singh et al.,, 2004, p. 6)\@BBCP it is that the agent reached ${\mathbf{s}}_{t+1}$ immediately from state ${\mathbf{s}}_{t}$—which it has just done—by following the skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$. If the agent expected, without uncertainty, to transition directly from ${\mathbf{s}}_{t}$ to ${\mathbf{s}}_{t+1}$, then $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ is 1 and the reward is 0 (the transition is very “unsurprising”). If the agent did not expect it could transition from ${\mathbf{s}}_{t}$ to ${\mathbf{s}}_{t+1}$ by following skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$, then $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ is 0 and the reward is 1 (maximally “surprising”). If the agent is uncertain, $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ will be somewhere in between. However, if the agent expected that it could reach ${\mathbf{s}}_{t+1}$ from ${\mathbf{s}}_{t}$ by following skill ${\mathbf{g}}_{{\mathbf{s}}_{t+1}}$, just that it would take more than a single step, $P^{{\mathbf{g}}_{{\mathbf{s}}_{t+1}}}$ will be discounted based on the length of the path. This reward encourages the agent to decrease the number of steps it takes to get from one state to another, strengthening its skill of reaching a particular state.
@@
Another example of C4.b, that is, being motivated by growing the number of skills in the agent’s repertoire, is the aforementioned VIC (see C2.a). The VIC reward function motivates the agent to grow its number of effective skills: in expectation, the second term of the reward function (Eq. 2, $\beta$) increases as the distribution over skills becomes more uniform. In theory, this encourages the distribution over skills not to collapse to a small number of skills,555\@BBOPcitet\@BAP\@BBNEysenbach et al., (2019, Appendix E, pp. 17–18)\@BBCP showed empirically that, in practice, since there are numerous other factors that affect how agents learn, this collapse may still happen. maintaining or increasing the size of the repertoire.
Advancing Competence Modelling in SDT
What is gained for human motivation researchers by identifying computational formalisms that match aspects of verbal theory? For one, translating a concept like competence into an algorithmic description induces a different way of thinking about theory: the process forces us to specify abstract and informal verbal theories to the point that they can be turned into a system that can actually produce observable behaviour. Observing how a computational model of a theory behaves can “reveal implicit assumptions and hidden complexities” \@BBOPcitep\@BAP\@BBN(Marsella et al.,, 2010, p. 4)\@BBCP and resolve jingle-jangle fallacies (\@BBOPcitep\@BAP\@BBNMurayama and Jach,, 2025, pp. 3–4, 6\@BBCP; \@BBOPcitep\@BAP\@BBNPekrun,, 2024, p. 26\@BBCP).
@@
For instance, LP-based methods can simultaneously exemplify skill use as (C2.a) an opportunity to use a particular skill, (C3.b) task performance as requiring a certain skill or skill level, and (C4.a) capacity growth as observing gains in the strength of one’s skills. Given that SDT does not acknowledge distinct facets, finding such a formalism that fits multiple facets supports the theory. On the other hand, in our matching, some facets are formally distinct. For example, it is unclear how (C1) effectance, (C4.b) being motivated by extending one’s repertoire of skills, or (C3.a) being motivated by observing that one performs well at an intended task, equate to each other and other facets. By identifying formalisms for each of the facets, and testing whether they result in demonstrably different observable behaviours (e.g., using simulation), we can prompt review of SDT—is one term, “competence”, being used for multiple different constructs?
@@
Another benefit of this matching is that it reveals underlying preconditions that SDT fails to specify, such as that a competence-motivated agent must be able to recognise when its actions cause effects (C1) and differentiate between distinct skills (C2, C3, C4). Similarly, formalisms of concepts such as novelty, diversity, and progress towards goals, underlie competence-driven learning in computational IM, yet are often overlooked in SDT. By drawing on such formalisms from computational agents that demonstrably and autonomously develop capacities in diverse environments, this paper shines light on overlooked preconditions and what potential mechanisms might computationally fulfil them.
We do not argue that identifying AI algorithms that demonstrably perform some capacities will mean that we have identified which algorithms best approximate human cognition \@BBOPcitep\@BAP\@BBN(Van Rooij et al.,, 2024, p. 626)\@BBCP. Rather, our work can support a cycle of theory development by inspiring new computational models, which can then be tested to refine the theory.
@@
Showing that RL formalisms can be aligned with SDT uncovers a wealth of resources for motivation research, with existing implementations of models, simulators, and evaluation metrics in the papers associated with each formalism. These resources can, by enabling large-scale simulation studies, allow us to explore computational IMs as candidates for model fitting to human data. We have used the different facets of competence outlined by \@BBOPcitet\@BAP\@BBNDeterding et al., (2024)\@BBCP as starting points for selecting promising computational formalisms. However, it remains an open question whether the facets capture different cognitive mechanisms or even require separate formalisms to generate associated explananda, such as exploratory and challenge-seeking behaviour. Towards investigating part of this question, the computational RL literature offers different methods for composing reward functions \@BBOPcitep\@BAP\@BBN(e.g., Bagot et al.,, 2020)\@BBCP, but how best to combine rewards and what methods would be psychologically plausible still require work. Such methods, in combination with the aforementioned resources, allow us to study how distinct facets of competence, either individually or in conjunction, contribute to competence-driven behaviour and motivation more widely.
These avenues of research are contingent on developing effective means to compare artificial agents with biological agents (a project partially undertaken by researchers in representational alignment; e.g., \@BBOPcitep\@BAP\@BBNSucholutsky et al.,, 2024\@BBCP). While work exists comparing behaviour induced by different intrinsic reward functions (e.g., \@BBOPcitep\@BAP\@BBNSantucci et al.,, 2012; Biehl et al.,, 2018; Linke et al.,, 2020\@BBCP), as does work studying different choices of state representation (e.g., \@BBOPcitep\@BAP\@BBNBurda et al.,, 2019\@BBCP), more research is needed on how different models behave. It is also an open question how to best evaluate intrinsically motivated open-ended learning \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022, pp. 1167–1169)\@BBCP.
@@
Most formalisms we identified relate to pursuing goals, so we need to understand how humans represent and generate goals \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2024, p. 1)\@BBCP. The topic has not been well studied for various reasons, e.g., computational intractability \@BBOPcitep\@BAP\@BBN(Byers et al.,, 2024, p. 1)\@BBCP. Excitingly, goals have recently emerged as a multidisciplinary area of research spanning AI, cognitive science, and developmental psychology (e.g., \@BBOPcitep\@BAP\@BBNColas et al.,, 2022; Chu et al.,, 2024; Davidson et al.,, 2024\@BBCP). \@BBOPcitet\@BAP\@BBNMolinaro and Collins, (2023, p. 1150)\@BBCP suggest that goals affect many agent dynamics, such as representing the environment, choosing relevant actions, and evaluating rewards.
@@
In summary, we demonstrate that computational models of IM can plausibly match verbal definitions of competence in SDT. Specifically, we show that four distinct facets of competence—effectance, skill use, task performance, and capacity growth—can be aligned with existing formalisms from intrinsically motivated RL, and in the process, highlight the potential of computational models to provide deeper insights into motivated behaviour. While this is a promising first step, further research is required to empirically study these computational models and consider other potential models, inviting collaboration from motivation researchers across disciplines.